The square of a positive number decreased by 4 times the number is 12. [only an algebraic soultion]

Answer by ewatrrr(24785)

 
Hi,
n^2 - 4n = 12
n^2 - 4n - 12 = 0 Factor
(n+2)(n-6) = 0 Using positive root only
n = 6
CHECKING our answer***
36 - 24 = 12
Wish You the Best in your Studies.